RC Update and Run plan for Sep 2,2020
RC Meeting Information
RC Daily Meetings are at 4:15 PM, remote
- Update on target work:
- polhe3, reference cell and the target ladder with new carbon hole foils installation is complete.
- Reference cell RTD & cable connection is complete and alignment is checked. https://logbooks.jlab.org/entry/3839245
- reference cell filled with H2 https://logbooks.jlab.org/entry/3839252 https://logbooks.jlab.org/entry/3839449
- Laser controlled access around 11pm yesterday to check the laser alignment, laser power and polarization. https://logbooks.jlab.org/entry/3839539 https://logbooks.jlab.org/entry/3839544 https://logbooks.jlab.org/entry/3839542
- the transverse right QWP that wasn't working earlier in the morning fixed by Ethan https://logbooks.jlab.org/entry/3839388 https://logbooks.jlab.org/entry/3839530
- Input connector of the EPR RF amplifier was broken. It was replaced by Chuck https://logbooks.jlab.org/entry/3839512
- EPR photo-diode replaced EPR Photodiode Replaced
- EPR calibration around 10-11pm today.
- ~ 3pm target enclosure is closed up. We're ready for beam
- Arun has the updated target encoder positions but cross checking with Bill before sending to Wei wei.
- Already check if the FSD clears out for all targets except for polHe3.
- FR coil is back on for 5-pass running. FR is off right now and once the beam is back it has to be initialized for the 5-pass beam energy. https://logbooks.jlab.org/entry/3839490
- 1-pass running has been scheduled for September 17/18.
- Accelerator has been working on the RF recovery since 10 am and expects to send beam around 6 pm.
- 5-pass running will start with reference cell running at low Q2 kinematics.
- Next Moeller during 5-pass running? (Note that Hall A will be doing Moeller on Friday)
- Notes for 1-pass:
- No Moeller measurement.
- Ask for beam energy measurement just after the pass change is complete.
General Reminders for Running
- Make sure the raster size is correct (4--4.5mm)
- Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
- Running on cell targets: raster always ON (4--4.5mm, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
- Always take 120Hz of EDTM
- Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
- Change the IHWP at least every 8 hours when running on glass cells. Call Hall A in advance to request. If there are any issues, call the RC
- Note that the autoload of IHWP status in the Start-Run GUI doesn't work anymore so please double check IHWP setting before you accept the GUI value.
Notices in Effect for Shift Crew
- Follow all COVID rules and regulations as well as remind each other. Additional food guild lines:
- All food are removed from the round table. Fruit are now store in the fridge, and individually packaged snacks are store on the shelf below the coffee machine.
- No food near the SL and TO control station. If you would like to eat something, please use the empty round table (used to store food in the past), that is now empty and sanitized.
- New coffee guide line: it is much preferred if you can bring you mug for the coffee. During the shift if one need to drink coffee near the control station, then the coffee mug must have a lid or have some kind of cover.
- Beam positions: 
- Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
- Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
- Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
- Ramp rate lock instructions are here and added to checklist
- An new reference Cell refill procedure: New procedure
- Magnet policies
- Spectrometer GUI (
go_magnets) should always be visible on Shift Leader computer, hcdesk3
- Slow ramping procedure of the SHMS, go small steps to the MOL, an great example (1 hour 30 min process)
- HMS Q1 must be ramped down in 50A steps in all cases
- Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
- New updates to the magnet momentum ramping procedure: https://hallcweb.jlab.org/wiki/index.php/How_to_operate_and_monitor_the_Spectrometer_Magnets
- If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog
- Spectrometer GUI (
- When rotating the spectrometer, use HCCAM01 to keep an eye on the SOLID Cherenkov test stand when rotating SHMS (in addition to watching the live angle-cam image, etc). The updated rotation instruction can be found: https://hallcweb.jlab.org/wiki/index.php/Spectrometer_Rotation#NOTICES_IN_EFFECT
- New SHMS matrix element
Planned Accelerator Configuration Changes
- Hall C to 1 pass on September 17/18, Thursday/Friday
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-126.96.36.199 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920